OpenStudy (anonymous):

multiply and simplify (x^1/3 +x^-1/3)^2

OpenStudy (anonymous):

use \[(a+b)^2=a^2+2ab+b^2\] with \[a=x^{\frac{1}{3}}\] and \[b=x^{-\frac{1}{3}}\]

OpenStudy (anonymous):

get \[x^{\frac{2}{3}}+2x^0+x^{-\frac{2}{3}}\] \[=x^{\frac{2}{3}}+2+x^{-\frac{2}{3}}\]